The RBCAB Corporation just paid a dividend of $1.13 per share. The company's CFO expects that the dividend will remain at that level for three years. After year three, it is expected that the dividend will grow at a rate of 4% indefinitely. If the required return is 10%, what is the value of stock today? At 5 years? At 10 years?
Answer #1

Current Dividend = D0 = $ 1.13, As the dividend remains constant for 3 Years, then D1 = D2 = D3 = $ 1.13

Required Return = 10 %, Growth Rate post Year 3 = 4 %

Current Stock Price = [1.13/(0.1-0.04)] x [1/(1.1)^(2)] = $ 15.5647 ~ $ 15.56

D3 = $ 1.13

Dividend at the end of Year 6 = D6 = 1.13 x (1.04)^(3) = $ 1.50403

Stock Price after Year 5 = D6 / (0.1-0.04) = 1.50403 / (0.1-0.04) = $ 25.067 ~ $ 25.07

Dividend at the end of Year 11 = D11 = 1.13 x (1.04)^(8) = $ 1.5465

Stock Price after Year 10 = D11/(0.1-0.04) = 1.5465 / (0.1-0.04) = $ 25.7747 ~ $ 25.77